OpenStudy (anonymous):

Find the largest integer value for the upper limit k if the two series are equal...

OpenStudy (anonymous):

\[\sum_{n}^{k}=1(n ^{2}+n)=\sum_{n}^{k}=1[(n-1)(n-2)(n-3)(n-4)(n-5)+n ^{2}+n]\]

OpenStudy (anonymous):

you think you could help me out with this problem?

OpenStudy (anonymous):

any tips, biddy?

OpenStudy (anonymous):

yeah, blacklabel :)

OpenStudy (anonymous):

any tips or hints?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

I just don't really know where to start with this problem :(

OpenStudy (anonymous):

do you think anyone else could help me?

OpenStudy (anonymous):

hello saeddiscover, do you think you could help me out with this problem?

OpenStudy (anonymous):

That sounds interesting! I will try ...

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

what is lower limit of n ?? It is needed to evaluate the summation...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

would the lower limit be 1?

OpenStudy (anonymous):

It may be. Are you given any lower limit for n? it should be known ..

OpenStudy (anonymous):

I'm not given it

OpenStudy (anonymous):

unless the problem has a typo or something

OpenStudy (anonymous):

supposing it was 1...

OpenStudy (anonymous):

the lower limit?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

how would I write it out as?

OpenStudy (anonymous):

I think it is an unwise question, it has no solution.

OpenStudy (anonymous):

but if the lower limit was 1, how would I solve it?

OpenStudy (anonymous):

what is equal sign (=) for in right hand side of your equation?

OpenStudy (anonymous):

my professor wrote the problem during the weekend; do you think he made a typo?

OpenStudy (anonymous):

Oops! I meant left-hand side of your equation.

OpenStudy (anonymous):

phi!!!!!! do you think you could help me out with this problem :)?

OpenStudy (phi):

out of curiosity, are you sure that is a one in 1(n^2 + n) people don't put in leading ones... so maybe it's a typo or means something else ?

OpenStudy (phi):

oh, I think it is part of n=1 (the lower limit) \[\sum_{n=1}^{k}(n^2+n)\]

OpenStudy (anonymous):

so that is wat the lower limit is?

OpenStudy (anonymous):

my bad

OpenStudy (phi):

yes, the =1 part belongs below the sigma (summation sign)

OpenStudy (anonymous):

so, what do I do next? The reason I have so much trouble with these problems is because I have a Professor that has trouble speaking English and trouble expressing his ideas :(.

OpenStudy (anonymous):

so anymore tips, phi?

OpenStudy (phi):

if we write it like this \[ \sum_{n=1}^{k}(n^2+n)= \sum_{n=1}^{k}(n^2+n) + \sum_{n=1}^{k}(n-1)(n-2)(n-3)(n-4)(n-5)\] it seems the complicated sum must be zero : \[ \sum_{n=1}^{k}(n-1)(n-2)(n-3)(n-4)(n-5)=0\]

OpenStudy (phi):

that suggests that as you increase n from 1 to 2 to 3, you keep getting 0 until n is bigger than 5 so we can go from n=1 to n=5 and get a sum of zero

OpenStudy (anonymous):

wait so

OpenStudy (anonymous):

(1-1)(2-2)(3-3)(4-4)(5-5)?

OpenStudy (anonymous):

But there are no (n-6)(n-7)(n-8)...

OpenStudy (anonymous):

so is the upper limit 5?

OpenStudy (anonymous):

the professor seems to be very crazy...

OpenStudy (phi):

the sum \[ \sum_{n=1}^{k}(n-1)(n-2)(n-3)(n-4)(n-5) \] means, add up (1-1)(1-2)(1-3)(1-4)(1-5) + (2-1)(2-2)(2-3)(2-4)(2-5) + and so on

OpenStudy (anonymous):

phi, sorry about earlier, apparently open study had to reboot

OpenStudy (phi):

notice that the first term in the summation is (1-1)(1-2)(1-3)(1-4)(1-5) and (1-1) is 0, so this turns into 0 the second term (2-1)(2-2)(2-3)(2-4)(2-5) is also 0 (because you have (2-2) in it) each term will be zero... when you get to n=5 you have the term (5-1)(5-2)(5-3)(5-4)(5-5) and again you get 0 (because you are multiplying by (5-5)=0 if you go to n=6, that term is (6-1)(6-2)(6-3)(6-4)(6-5) and this is NOT zero. so the biggest k can be (and still get a sum of zero) is k=5

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

incredible

OpenStudy (anonymous):

so phi

OpenStudy (anonymous):

is that the solution to the problem?

OpenStudy (anonymous):

how bout the n squared+n part?

OpenStudy (phi):

to summarize \[ \sum_{n=1}^{k}(n^2+n)= \sum_{n=1}^{k}(n^2+n) +(n-1)(n-2)(n-3)(n-4)(n-5) \\ \sum_{n=1}^{k}(n^2+n)= \sum_{n=1}^{k}(n^2+n) + \sum_{n=1}^{k}(n-1)(n-2)(n-3)(n-4)(n-5) \] subtract \(\sum_{n=1}^{k}(n^2+n)\) from both sides to get \[ 0=\sum_{n=1}^{k}(n-1)(n-2)(n-3)(n-4)(n-5) \] you can go up to n=5 and get a sum of zero \[ (1-1)(1-2)(1-3)(1-4)(1-5) + (2-1)(2-2)(2-3)(2-4)(2-5) +...+\\ (5-1)(5-2)(5-3)(5-4)(5-5) =0\]

OpenStudy (phi):

if you go to n=6, that term is (6-1)(6-2)(6-3)(6-4)(6-5) and this is NOT zero. so the biggest k can be (and still get a sum of zero) is k=5

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

you are such a big help

OpenStudy (anonymous):

I give you an S+